OpenStudy (anonymous):

Let f(x) = 3x2 + x - 3 and g(x) = x2 - 5x + 1. Find f(x) - g(x). 2x2 - 4x - 2 2x2 - 4x - 4 2x2 + 6x - 2 2x2 + 6x - 4

OpenStudy (anonymous):

Let’s re-write the function as follows: y = (x2 + x)5(1 + x3)2 Let’s consider a function k as follows: k = x2 + x k' = 2x + 1 Let’s consider a function h as follows: u = k5 = (x2 + x)5 u' = nkn-1k' = 5(x2 + x)4(2x + 1) = 5(2x + 1)(x2 + x)4 Let’s consider a function g as follows: g = 1 + x3 g’ = 3x2 Let’s consider a function m as follows: v = g2 = (1 + x3)2 v' = ngn-1g' = 2(1 + x3)(3x2) = 6x2(1 + x3) Let’s consider a function f as follows: f = uv To calculate the derivative, we use the formula (uv)' = u'v + uv': f' = u'v + uv' = 5(2x + 1)(x2 + x)4(1 + x3)2 +(x2 + x)56x2(1 + x3) = 5(2x + 1)(x2 + x)4(1 + x3)2 +6x2(1 + x3)(x2 + x)5 = (x2 + x)4(1 + x3)[5(2x + 1)(1 + x3) + 6x2(x2 + x)]